OpenStudy (anonymous):

If 0.765g of an unknown gas were collected at 741 torr and 95C in a 245 mL container, then what is the molecular weight of this unknown gas?

OpenStudy (anonymous):

just use: pV= nRT from there calculate "n" and then use: n= m/M -> M=m/n

OpenStudy (anonymous):

just be careful about units, temperature in K, pressure in Pa, volume in m3, R=8,314 JK-1mol-1